(7y=4t^2)dy+4tydt=0


1
Expert's answer
2021-06-10T02:48:48-0400
"(7y+4t^2)dy+4tydt=0"

"P(t, y)=7y+4t^2"

"Q(t, y)=4ty"

"\dfrac{\partial P}{\partial t}=8t"

"\dfrac{\partial Q}{\partial y}=4t"

"\dfrac{\partial P}{\partial t}\not=\dfrac{\partial Q}{\partial y}"

Let


"y(7y+4t^2)dy+4ty^2dt=0"

"M(t, y)=7y^2+4t^2y"

"N(t, y)=4ty^2"

"\dfrac{\partial M}{\partial t}=8ty"

"\dfrac{\partial N}{\partial y}=8ty"

"\dfrac{\partial M}{\partial t}=\dfrac{\partial N}{\partial y}"

"\dfrac{\partial F}{\partial y}=7y^2+4t^2y"

"F(y, t)=\dfrac{7}{3}y^3+2t^2y^2+\psi(t)=c_1"

"\dfrac{\partial F}{\partial t}=4ty^2+\dfrac{d\psi}{dt}=4ty^2=>\dfrac{d\psi}{dt}=0=>\psi=c_2"



Then


"\dfrac{7}{3}y^3+2t^2y^2=c"